OpenStudy (anonymous):

can 7/50 be simplified??

OpenStudy (ryan123345):

no

OpenStudy (anonymous):

kk thx

OpenStudy (ryan123345):

welcome :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

0.14

OpenStudy (anonymous):

30/100 can be simplified what??

OpenStudy (anonymous):

0.3

OpenStudy (anonymous):

by 3??

OpenStudy (anonymous):

30/100 =0.3

OpenStudy (ryan123345):

no. @lovemenplayhard he is talking if it can be simplified not divided lol

OpenStudy (ryan123345):

and yes it can be simplified. just divide both numbers by 10 :)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

division is also simplification and there is no limit of simplification LOL and 30/100 = 3/10

OpenStudy (anonymous):

can 30/100 be simplified by 10??

OpenStudy (ryan123345):

yes

OpenStudy (anonymous):

i got 3/10

OpenStudy (ryan123345):

and thats when stop, You cant simplify it any more.

OpenStudy (anonymous):

im learning about about percents. when u have 5% do u put it over 10 or 100??

OpenStudy (anonymous):

@ryan123345

OpenStudy (ryan123345):

well it depends. Are you learning about.

OpenStudy (ryan123345):